# SpIDER: Spatially Informed Dense Embedding Retrieval for Software Issue Localization

### v1

@@ 6.2 Results
Java SweRankEmbed-Small ZS SpIDER 0.37 0.28 0.08
DER 0.32 0.24 0.08
CodeSAGE-Small ZS SpIDER 0.21 0.19 0.05
DER 0.15 0.13 0.03
CodeSAGE-Small-PEFT SpIDER 0.31 0.25 0.10
DER 0.23 0.17 0.09
BM25 SpISR 0.23 0.17 0.07
SR 0.22 0.17 0.05
JavaScript SweRankEmbed-Small ZS SpIDER 0.39 0.30 0.18
DER 0.31 0.23 0.15
CodeSAGE-Small ZS SpIDER 0.17 0.12 0.08
DER 0.09 0.06 0.05
CodeSAGE-Small-PEFT SpIDER 0.32 0.24 0.13
DER 0.25 0.18 0.11
BM25 SpISR 0.21 0.14 0.08
SR 0.14 0.10 0.06
TypeScript SweRankEmbed-Small ZS SpIDER 0.52 0.42 0.08
DER 0.40 0.32 0.23
CodeSAGE-Small ZS SpIDER 0.33 0.26 0.04
DER 0.26 0.21 0.13
CodeSAGE-Small-PEFT SpIDER 0.30 0.26 0.05
DER 0.19 0.15 0.12
BM25 SpISR 0.32 0.27 0.04
SR 0.24 0.21 0.11
SWEBench-Verified
Python SweRankEmbed-Small ZS SpIDER 0.61 0.56 0.32
DER 0.54 0.49 0.29
CodeSAGE-Small ZS SpIDER 0.30 0.27 0.13
DER 0.21 0.19 0.10
CodeSAGE-Small-PEFT SpIDER 0.55 0.49 0.27
DER 0.43 0.38 0.22
BM25 SpISR 0.45 0.41 0.25
SR 0.38 0.34 0.22
Improvement is Consistent Across Datasets, Metrics, % Languages
Next, we augment all baseline retrieval methods—including dense models (SweRankEmbed-Small, codesage-small-v2) and BM25 sparse retrievers– with code graph structure information and measure the resulting improvements. We perform this analysis across all datasets (SWEBench-Verified, SWE-PolyBench, Multi-SWEBench) as reported in Tab. 2 and 3.
Our results reveal that our SpIDER strategy consistently enhances retrieval performance, particularly in Recall@K and Acc@K, across all programming languages and datasets. The best-performing embedding model across datasets is SweRankEmbed-Small, and applying SpIDER further increases Recall@20 and Acc@20 by at least $12.9\%$ and $14\%$, respectively.
Fig. 6 shows that SpIDER consistently outperforms baseline DER methods across varying numbers of function-level edits when using SweRankEmbed-Small. With the LLM temperature set to 0.1, the standard deviation across multiple iterations was $\leq 0.01$ for all metrics, indicating stable performance.
MRR@$K$ Gains Depend on Graph Neighborhoods
Retrieval Strategy is Agnostic to Downstream Rerankers
Our retrieval strategy is agnostic to downstream rerankers or localization steps. To validate this, we evaluate an LLM selector (similar to the reranker in reddy2025sweranksoftwareissuelocalization) that retrieves the top 3 modules from a set of $K=20$ candidates. As reported in Tab. 2, improving top-$K$ coverage with SpIDER also translates to notable gains in Recall@20 and Acc@20 for top-3 localized modules.
@@ 6.3 Ablation Experiments
$\boldsymbol{C}$ Python JavaScript Java TypeScript
1 0.47 0.38 0.29 0.56 0.48 0.18 0.32 0.24 0.18 0.32 0.24 0.06
3 0.49 0.39 0.30 0.59 0.51 0.18 0.35 0.26 0.19 0.31 0.24 0.06
5 0.49 0.40 0.30 0.60 0.52 0.18 0.36 0.27 0.18 0.35 0.28 0.06
7 0.49 0.39 0.30 0.60 0.52 0.18 0.36 0.27 0.19 0.37 0.30 0.06
9 0.48 0.39 0.30 0.60 0.53 0.19 0.37 0.27 0.19 0.39 0.31 0.06
Table 5: Impact of the number of neighbors per center ($d$) on retrieval performance and LLM token usage using SpIDER + SweRankEmbed-Small for $K=20$, $N=100$, $C=3$. Increasing $d$ improves Recall@20 and Accuracy@20 across languages, while LLM input/output tokens scale with $d$. Best values per metric are highlighted.
$\boldsymbol{d}$ Python JavaScript Java TypeScript
Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output
2 0.45 0.36 2,642 32 0.54 0.46 2,032 50 0.33 0.25 2,158 22 0.31 0.23 1,474 21
4 0.46 0.37 3,591 35 0.57 0.49 3,115 25 0.33 0.25 2,083 54 0.30 0.24 1,994 24
6 0.50 0.41 11,904 48 0.62 0.53 10,065 42 0.33 0.25 3,615 67 0.36 0.28 4,334 33
8 0.55 0.44 29,658 69 0.67 0.58 15,028 48 0.34 0.27 5,594 82 0.36 0.28 5,643 37
On the Number of Top $C$ Centers
In this section, we provide ablation studies on varying the number of top $C$ centers on SWE-PolyBench for retrieval performance. Increasing $C$ results in more neighborhood explorations and addition of relevant neighbors in the list of top-$K$ retrieved modules at the cost of dropping bottom semantically ranked modules. Furthermore, each neighborhood exploration around a center requires 1 LLM call for filtering neighbors. Hence, increasing $C$ results in increased number of LLM calls. In table˜4, we ablate on $C$ and observe the corresponding performance changes. We see that the performance saturates at $C=5$ for $N=500,d=4$ at $K=20$. Hence, we stick with $C=5$ for all our experiments. A downstream user can aim to choose a $C$ that fits in their LLM invocation limits per instance.
On $K$
The choice of $K$ for retrieval depends on the budget of the downstream application. Retrieval performance is expected to improve as $K$ increases. We further validate this claim in figures 2 and 5 where we see a monotonic improvement in performance as $K$ incremented. The downstream user can choose a $K$ that is allowed by their budget while aiming for best retrieval performance.
On depth $d$
In this section, we vary the maximum depth $d$ for breadth-first search (BFS) exploration while selecting neighbors and observe the corresponding impact on the performance. Greater depth would result in more number of neighbors per center that need to be filtered using the two-stage neighborhood filtering. In table˜5, we ablate on $d$ keeping $N=100,C=3$ and $K=20$. We see that there is notable improvement in the Recall@20 and Acc@20 performance with increasing depth $d$, however, increasing $d$ also increases the token consumption to LLM while filtering the neighbors. The downstream user can choose $d$ based on their token budget per LLM invocation.
On the Primary Neighborhood Filtering Threshold NN
In table˜12, ablate on $N$ to see it’s impact on the retrieval performance. Higher $N$ results in more neighbors passing the primary filteration and hence, more neighbors for the LLMs to sift from. We see that the performance saturates around $N=300$ for $C=3,K=20,d=4$. Hence, the choice of $N$ offers the same flexibility as $d$ to the downstream user. The user can choose $N$ based on the token limit per LLM call.
@@ Appendix A Appendix
A.1 Graph construction details
@@
A.2 Extension to Class-level edits
@@
TypeScript SweRankEmbed-Small ZS SpIDER
DER
CodeSAGE-Small ZS SpIDER
DER
CodeSAGE-Small-PEFT SpIDER
DER
BM25 SpISR
SR
In table˜6, we demonstrate the efficacy of the proposed approach at a class level. Here, we include only Python and Java here, excluding JavaScript, as most JavaScript repositories in the SWE-PolyBench dataset lack significant use of class syntax (introduced in ECMAScript 2015), resulting in very few instances with class ground truth nodes. A ground truth edit is classified as a class-level edit (and the corresponding class as a ground truth class node), if the patch introduces a modification within any part of the class declaration or definition (e.g., constructors, destructors, or methods) to address the issue description.
A.3 Additional Experiments

### v2

@@ 5.2 Results
Java SweRankEmbed-Small ZS SpIDER 0.36 0.27 0.25 0.17 0.37 0.28
DER 0.31 0.24 0.24 0.16 0.32 0.24
CodeSAGE-Small ZS SpIDER 0.23 0.17 0.16 0.09 0.21 0.19
DER 0.15 0.10 0.11 0.07 0.15 0.13
CodeSAGE-Small-PEFT SpIDER 0.28 0.18 0.20 0.14 0.31 0.25
DER 0.21 0.14 0.17 0.11 0.23 0.17
BM25 SpISR 0.16 0.10 0.12 0.08 0.23 0.17
SR 0.15 0.10 0.12 0.09 0.22 0.17
JavaScript SweRankEmbed-Small ZS SpIDER 0.60 0.52 0.43 0.37 0.39 0.30
DER 0.52 0.45 0.41 0.34 0.31 0.23
CodeSAGE-Small ZS SpIDER 0.38 0.32 0.30 0.25 0.17 0.12
DER 0.26 0.21 0.22 0.17 0.09 0.06
CodeSAGE-Small-PEFT SpIDER 0.55 0.48 0.41 0.35 0.32 0.24
DER 0.45 0.38 0.36 0.30 0.25 0.18
BM25 SpISR 0.46 0.39 0.36 0.30 0.21 0.14
SR 0.36 0.29 0.31 0.25 0.14 0.10
TypeScript SweRankEmbed-Small ZS SpIDER 0.35 0.28 0.27 0.21 0.52 0.42
DER 0.29 0.21 0.25 0.19 0.40 0.32
CodeSAGE-Small ZS SpIDER 0.16 0.10 0.14 0.09 0.33 0.26
DER 0.10 0.06 0.09 0.06 0.26 0.21
CodeSAGE-Small-PEFT SpIDER 0.25 0.18 0.22 0.16 0.30 0.26
DER 0.19 0.13 0.18 0.12 0.19 0.15
BM25 SpISR 0.11 0.09 0.10 0.07 0.32 0.27
SR 0.07 0.05 0.07 0.05 0.24 0.21
Table 3: Ablation on number of seed centers $C$ using SweRankEmbed-Small+SpIDER. Increasing $C$ improves Recall@20 and Accuracy@20 across languages up to saturation.
$\boldsymbol{C}$ Python JavaScript Java TypeScript
Recall@20 Acc@20 Recall@20 Acc@20 Recall@20 Acc@20 Recall@20 Acc@20
1 0.47 0.38 0.56 0.48 0.32 0.24 0.32 0.24
3 0.49 0.39 0.59 0.51 0.35 0.26 0.31 0.24
5 0.49 0.40 0.60 0.52 0.36 0.27 0.35 0.28
7 0.49 0.39 0.60 0.52 0.36 0.27 0.37 0.30
9 0.48 0.39 0.60 0.53 0.37 0.27 0.39 0.31
Table 4: Ablation on BFS exploration depth ($d$) using SpIDER + SweRankEmbed-Small for $K=20$, $N=100$, $C=3$. Increasing $d$ improves Recall@20 and Accuracy@20 across languages, while LLM input/output tokens scale with $d$.
$\boldsymbol{d}$ Python JavaScript Java TypeScript
Recall@20 Acc@20 Tokens Recall@20 Acc@20 Tokens Recall@20 Acc@20 Tokens Recall@20 Acc@20 Tokens
Input Output Input Output Input Output Input Output
2 0.45 0.36 2.6K 32 0.54 0.46 2.0K 50 0.33 0.25 2.2K 22 0.31 0.23 1.5K 21
4 0.46 0.37 3.6K 35 0.57 0.49 3.1K 25 0.33 0.25 2.1K 54 0.30 0.24 2.0K 24
6 0.50 0.41 11.9K 48 0.62 0.53 10.1K 42 0.33 0.25 3.6K 67 0.36 0.28 4.3K 33
8 0.55 0.44 29.7K 69 0.67 0.58 15.0K 48 0.34 0.27 5.6K 82 0.36 0.28 5.6K 37
Leveraging code graphs improves retrievaltable˜2 demonstrates the benefit of incorporating code graph structure into retrieval. Comparing standard dense (DER) and sparse (SR) retrieval with our graph-aware variant (SpIDER) across Python, TypeScript, JavaScript, and Java repositories from SWE-PolyBench, Multi-SWEBench, and SWEBench-Verified at $K=20$, we observe consistent improvements across all datasets and languages.
Fig. 2 further illustrates this trend on SWE-PolyBench as the retrieval budget $K$ varies. SweRankEmbed-Small + SpIDER consistently outperforms all baselines across languages and $K$ values, highlighting the effectiveness of combining semantic similarity with explicit modeling of spatial relationships and LLM-based neighborhood filtering. Notably, the zero-shot SweRankEmbed-Small model—trained exclusively on roughly 3,300 Python repositories—remains competitive on non-Python codebases, outperforming BM25 and codesage-small-v2 in both zero-shot and PEFT settings, demonstrating strong cross-language transfer.
Across datasets, SpIDER improves Recall@20 and Acc@20 of SweRankEmbed-Small by at least $13\%$ and $14\%$, respectively, indicating that graph-aware retrieval yields substantial performance gains even when paired with already strong semantic embeddings.
Improvement is consistent across datasets, metrics, and languages  As shown in fig.˜3, SpIDER yields consistent improvements across both context configurations, achieving gains of +0.46+0.46 percentage points (pp) at $K=5$ and +2.74+2.74 pp at $K=10$, corresponding to 2 and 12 additional resolved instances, respectively. These results demonstrate that enhanced retrieval coverage directly translates into measurable gains in downstream code generation performance, completing the chain from graph-aware retrieval to end-task success.
Retrieval strategy is agnostic to downstream rerankers  Our retrieval strategy is designed to operate independently of downstream reranking or localization components. To validate this property, we evaluate all retrieval methods in conjunction with an LLM-based reranker following Reddy et al. (2025), which selects the top-3 functions from $K=20$ retrieved candidates based on its reasoning capabilities.
As shown in table˜2, the improved top-20 coverage provided by SpIDER consistently translates into higher Recall@3 and Acc@3 after reranking. This implies that gains achieved at the retrieval stage are preserved and often amplified by downstream reranking, and that SpIDER is complementary to, rather than coupled with, specific localization strategies.
Improved retrieval translates to better code generation  To evaluate the practical impact of improved retrieval on downstream task performance, we pass the retrieved function paths to swe-mini-agent Yang et al. (2024) executed with Claude Sonnet 4.5, measuring pass rate on the SWE-Bench-Verified benchmark with retrieval budgets $K=5$ and $K=10$.
As shown in fig.˜3, SpIDER yields consistent improvements across both context configurations, achieving gains of +0.46+0.46 percentage points (pp) at $K=5$ and +2.74+2.74 pp at $K=10$, corresponding to 2 and 12 additional resolved instances, respectively. These results demonstrate that enhanced retrieval coverage directly translates into measurable gains in downstream code generation performance, completing the chain from graph-aware retrieval to end-task success.
Figure 3: Pass rate improvement of SWERankEmbed-Small-SpIDER over SWERank on SWE-bench Verified with $C=5,d=4,N=500$.
@@ 5.3 Ablation Experiments
To better understand how SpIDER’s hyperparameters influence retrieval performance and to provide actionable guidance for downstream users, we conduct ablation studies on four key parameters: number of seed centers $C$, retrieval budget $K$, BFS exploration depth $d$ and primary neighborhood filtering threshold $N$. For each hyperparameter, we follow a consistent structure: we first describe the motivation for the parameter, then its effect on retrieval performance, discuss trade-offs in terms of computational cost/LLM usage, and finally provide guidance for tuning in practice. These experiments aim to help users configure SpIDER effectively for different codebases and resource budgets.
Number of Top $C$ Centers  The number of seed centers $C$ determines how many top-ranked nodes are used as starting points for neighborhood exploration. Increasing $C$ expands the explored neighborhood, adding more relevant neighbors to the top-$K$ retrieved nodes but potentially displacing lower-ranked semantic candidates. Each neighborhood exploration requires one LLM call, so higher $C$ increases total LLM invocations.
table˜3 shows that performance saturates at $C=5$ for $N=500$, $d=4$, and $K=20$, indicating that exploring neighbors of the top-ranked nodes efficiently recovers ground-truth functions while keeping LLM usage reasonable. Users can adjust $C$ according to their LLM budget.
Retrieval Budget $K$  The retrieval budget $K$ controls the number of candidate nodes returned. As expected, performance improves monotonically with increasing $K$ (figs.˜2 and 6), since a larger pool increases the chance of including relevant nodes. Users can select $K$ based on the constraints of downstream tasks, balancing retrieval quality against processing cost.
Exploration Depth $d$  The maximum BFS depth $d$ determines how many neighbors are explored around each seed center. Larger $d$ increases the number of candidate neighbors, improving Recall@20 and Acc@20 (table˜4) but also increasing token consumption during the LLM-based neighborhood filtering. Users should select $d$ considering the trade-off between retrieval gains and LLM resource usage.
Primary Neighborhood Filtering Threshold NN  The threshold $N$ controls how many neighbors pass the initial semantic filter before LLM-based refinement. Higher $N$ allows more neighbors to be considered, increasing the chance of recovering relevant nodes but also consuming more LLM tokens. Ablations in table˜11 show performance saturates around $N=300$ for $C=3$, $d=4$, and $K=20$. Like $d$, NN can be tuned based on downstream LLM token budgets.
Takeaways  Overall, these ablations show that SpIDER’s hyperparameters offer a clear trade-off between retrieval performance and LLM cost. Key settings i.e., $C=5$, $d=4$, $N=500$, and $K=20$ provide strong performance across SWE-PolyBench, while users can tune $C$, $d$, NN, and $K$ according to resource constraints. Our results highlight that carefully leveraging top-ranked nodes and their neighborhoods is an effective strategy for maximizing retrieval coverage while controlling computational cost.
@@ Appendix A Appendix
A.1 Theoretical Analysis
The proof for proposition˜4.2 is simple. We start by providing a theoretical justification for why incorporating graph locality and LLM-based neighbor selection improves retrieval performance over standard dense embedding retrieval. Our analysis focuses on recall at a fixed budget $K$ and relies on mild structural and statistical assumptions.
A.2 Modeling Assumptions
We follow the problem definition from section˜3 and use $\mathcal{S}_{K}$ to denote $\mathcal{S}_{K}(Q)$ for notational convenience. Then we make the following assumptions.
Assumption 1 (Score Distributions).
There exist distributions $F_{1}$ and $F_{0}$ on $\mathbb{R}$ such that
$s_{Q}(v)\sim\begin{cases}F_{1},&v\in\mathcal{V}^{*},\\ F_{0},&v\notin\mathcal{V}^{*},\end{cases}$
where $F_{i}(t)=\Pr[s_{Q}(v)\leq t\mid s_{Q}(v)\sim F_{i}]$ and $F_{1}$ first-order stochastically dominates $F_{0}$, i.e.,
$F_{1}(t)\leq F_{0}(t)\quad\forall t\in\mathbb{R}.$
(See Shaked and Shanthikumar [2007] for background on stochastic dominance.)
Assumption 2 (Graph Locality of Relevance).
There exists an integer $d\geq 1$ such that for any $v\in\mathcal{V}^{*}$,
$\Pr\big(u\in\mathcal{V}^{*}\mid u\in\Gamma_{d}(\{v\})\big)\geq\Pr(u\in\mathcal{V}^{*}).$
That is, relevance is non-negatively correlated with graph proximity.
Assumption 3 (LLM Selector Accuracy).
For any $u\in\mathcal{C}_{Q}$,
$\Pr(\mathcal{L}(u)=1\mid u\in\mathcal{V}^{*})$ $=\alpha$
$\Pr(\mathcal{L}(u)=1\mid u\notin\mathcal{V}^{*})$ $=\beta,$
where $0\leq\beta<\alpha\leq 1$.
A.3 Expected Recall Change
Define
$B_{Q}$ $=\mathcal{S}_{K}\setminus\mathcal{S}_{K-|\mathcal{U}_{Q}|}$
$\pi_{B}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in B_{Q}),$
$\pi_{\Gamma}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in\hat{\mathcal{C}}_{Q}\setminus\mathcal{S}_{K}).$
Here, $B_{Q}$ refers to bottom $|\mathcal{U}_{Q}|$ ranked nodes in $\mathcal{S}_{K}(Q)$.
Proposition A.1 (Expected Recall Difference).
Under Assumptions 1–3, the expected change in recall satisfies
$\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}^{\texttt{LLM}}_{K})]-\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}_{K})]=\frac{\alpha\,\mathbb{E}[|\mathcal{U}_{Q}\cap\mathcal{V}^{*}|]-\mathbb{E}[|B_{Q}\cap\mathcal{V}^{*}|]-\beta\,\mathbb{E}[|\mathcal{U}_{Q}\setminus\mathcal{V}^{*}|]}{|\mathcal{V}^{*}|}.$
Proof.
By construction,
$|\mathcal{S}^{\texttt{LLM}}_{K}\cap\mathcal{V}^{*}|=|\mathcal{S}_{K}\cap\mathcal{V}^{*}|-|B_{Q}\cap\mathcal{V}^{*}|+|\mathcal{U}_{Q}\cap\mathcal{V}^{*}|.$
Taking expectations and dividing by $|\mathcal{V}^{*}|$ yields the result. ∎
A.4 Sufficient Condition for Improvement
Proposition A.2 (Restating Sufficient Condition for Recall Improvement).
Under the assumptions 1, 2 and 3 listed in section˜A.2, if
$\alpha\pi_{\Gamma}>\pi_{B}+\beta(1-\pi_{\Gamma})$, then it holds that
$\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}^{\texttt{LLM}}_{K})]>\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}_{K})].$
Proof.
By linearity of expectation,
$\mathbb{E}[|\mathcal{U}_{Q}\cap\mathcal{V}^{*}|]=\alpha K\pi_{\Gamma},\quad\mathbb{E}[|\mathcal{U}_{Q}\setminus\mathcal{V}^{*}|]=K(1-\pi_{\Gamma}).$
Similarly,
$\mathbb{E}[|B_{Q}\cap\mathcal{V}^{*}|]=K\pi_{B}.$
Substituting into Proposition A.1 and simplifying yields the stated claim. This concludes the proof for proposition˜4.2
A.5 Graph-Structured Relevance
We now show that $\pi_{\Gamma}$ is strictly larger than $\pi_{B}$ under a standard cluster assumption.
Definition (Relevant Subgraph).
Let $H_{q}=(V_{q},E_{q})$ be an induced subgraph such that $\mathcal{V}^{*}\subseteq V_{q}$ and
$\operatorname{diam}(H_{q})\leq d.$
Proposition A.3 (Graph Expansion Enriches Relevance).
Suppose:
  1. 1.
    There exists at least one seed $v\in C_{q}\cap V_{q}$.
  2. 2.
    Assumptions 1 and 2 hold.
Then
$\pi_{\Gamma}\geq\Pr(u\in\mathcal{V}^{*}),$
with strict inequality unless $\mathcal{V}^{*}=V_{q}$.
Proof.
Since $\operatorname{diam}(H_{q})\leq d$, all nodes in $V_{q}$ lie in $\Gamma_{d}(\{v\})$. By Assumption 2, conditioning on membership in $\Gamma_{d}(\{v\})$ increases the probability of relevance. Restricting further to $\operatorname{TopN}$ preserves this enrichment by Assumption 1 (stochastic dominance). ∎
A.6 Conclusion
The above results show that when (i) relevant functions form a localized region in the code graph, (ii) dense retrieval places at least one seed in that region, and (iii) the LLM selector has higher true-positive than false-positive rate, then graph-aware retrieval strictly improves expected recall at fixed budget $K$.
A.7 Algorithm
Algorithm 1 SpIDER
0: Code graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, issue description $Q$, bi-modal encoder $\mathcal{F}(\cdot)$
0: Parameters: retrieval budget $K$, filtering threshold NN, number of centers $C$, search depth $d$
1:// Semantic Retrieval
2: Compute node embeddings: $\{\mathcal{F}(v)\}_{v\in\mathcal{V}}$
3: Compute query embedding: $\mathcal{F}(Q)$
4: Compute relevance scores: $s_{Q}(v)=\cos(\mathcal{F}(v),\mathcal{F}(Q))$ for all $v\in\mathcal{V}$
[…]
7:// Seed Selection
[…]
9:// Neighborhood Exploration
[…]
11:// Two-Stage Neighbor Filtering
[…]
14:// Output Construction
[…]
16:return $\mathcal{S}(Q)$
Note: The LLM selector $\mathcal{L}$ receives the source code content of all candidates $u\in\hat{\mathcal{C}}_{Q}$ for a given center, along with the issue description $Q$. It returns $\mathcal{L}(u)=1$ for candidates deemed relevant to resolving the issue, and 0 otherwise.
A.8 Graph construction details
@@
A.9 Extension to Class-level edits
@@
In table˜5, we demonstrate the efficacy of the proposed approach at a class level. Here, we include only Python and Java here, excluding JavaScript, as most JavaScript repositories in the SWE-PolyBench dataset lack significant use of class syntax (introduced in ECMAScript 2015), resulting in very few instances with class ground truth nodes. A ground truth edit is classified as a class-level edit (and the corresponding class as a ground truth class node), if the patch introduces a modification within any part of the class declaration or definition (e.g., constructors, destructors, or methods) to address the issue description.
A.10 Additional Experiments